OpenStudy (anonymous):

The Johnsons need a new living room couch. The couch they like costs $2,700. The salesperson told the Johnsons that the couch will be on sale next month for 20% off. If the sales tax rate in their city is 7.5%, how much will the Johnsons save including sales tax, if they wait until the couch goes on sale?

OpenStudy (anonymous):

To find out the price of something with percent added or subtracted, multiply the original price by 1 + the percent converted to decimal or 1- the percent converted to decimal.

OpenStudy (anonymous):

Convert a percent to decimal by taking the percent and move the decimal two places to the left or multiplying it by .01.

OpenStudy (anonymous):

To find the difference between the two prices, calculate the non sale price and the sale price and just subtract one from the other. remember that the 20% off applies before you have to add in the 7.5% sales tax.

OpenStudy (amistre64):

Percent problems have 3 parts to them: (Percent) = (Part amount)/(Total amount) times 100. This can be written 3 ways to solve for any of the missing parts. Fs = 2700(20%) Fs = 2700(.2) Fs = 540. The Furniture on sale(Fs) is $540 off its original price. How much tax would that have been? T = 540(7.5%) T = 540(.075) T = 40.50 The tax is $40.50. Total savings is Fs + T: 540 + 40.50 = $580.50

OpenStudy (anonymous):

Let me know if this helps or if you're still stuck on something.

OpenStudy (anonymous):

thank you very much both of you(: i understand this way better but isnt this a porpotion (however you speell it) problem?

OpenStudy (amistre64):

proportions, fractions, and percents are all interrelated. They are like one big happy kentucky family :)